OpenStudy (anonymous):

The height of an equilateral rectangular prism decreases by two units. The new lateral area is less than the original by how much? two times the area of the base two less than the perimeter of the base two less than the the area of the base two times the perimeter of the base

OpenStudy (anonymous):

two less than the the area of the base

OpenStudy (anonymous):

^what i think

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

The lateral area is LSA = 2*(LH+WH) if the height decreases by 2, then H becomes H-2, meaning that we get LSA = 2*(L(H-2)+W(H-2)) LSA = 2*(LH-2L+WH-2W) LSA = 2*(LH+WH-2L-2W) LSA = 2*(LH+WH)+2*(-2L-2W) LSA = 2*(LH+WH)-2*(2L+2W) Notice how the first piece 2*(LH+WH) is the original lateral surface area. Also, notice how the new lateral surface area is the result of subtracting off double the perimeter So it's really "two times the perimeter of the base"

jimthompson5910 (jim_thompson5910):

So decreasing the height by 2 means that the new lateral surface area is less by 2 times the perimeter of the base.